A jet flies at a speed of 275 m/s. The pilot fires a missile forward off a mounting at a

speed of 700 m/s relative to the ground. The respective masses of the jet and the

missile are 5 000 kg and 50 kg. Calculate the speed of the jet immediately after the

missile had been fired.



1
Expert's answer
2021-10-25T15:38:28-0400

by impulse conservation law:

"v(m_1+m_2)=m_1v_1+m_2v_2"

where v is jet speed before the missile had been fired,

v1 is missile speed,

v2 is jet speed after the missile had been fired,

m1 and m2 are masses of missile and jet.


"v_2=\frac{v(m_1+m_2)-m_1v_1}{m_2}=\frac{275(5000+50)-700\cdot 50}{5000}=270.75" m/s